OpenStudy (anonymous):

Write an equation of the line containing the given point and perpendicular to the given line. (0,6); 9x+8y=5

OpenStudy (alexwee123):

first isolate y... 9x+8y=5 8y=-9x+5 y=(-9/8)*x+5/8 the perpendicular slope of -9/8 is 8/9 so sub that in with the given point into point slope equation y-y1=m(x-x1) so... y-6=8/9(x-0) so y=(8/9)*x+6 would be your answer